OpenStudy (anonymous):

f(5x)=x/x^2+1 what is f(x)........?

OpenStudy (anonymous):

\[\huge \frac{\frac{x}{5}}{(\frac{x}{5})^2 + 1}\]

OpenStudy (anonymous):

just simplify